For salts: 1) CrCl 3 ; 2) Na 2 CO 3 ; 3) NaNO 2 ; 4) KCl, compose the shortened ionic-molecular and molecular hydrolysis equations for stage I and determine the pH of the solution (<7,> 7 = 7, ≈7).


1
Expert's answer
2021-06-14T13:37:30-0400

1) CrCl3+H2O=CrOHCl2 + HCl

Cr(3+)+3Cl(1-)+H2O=CrOH(2+)+2Cl(2-)+H(1+)+Cl(1-)

Cr(3+)+H2O=CrOH(2+)+H(1+)

pH<7


2) Na2CO3 + H2O=NaHCO3 + NaOH

2Na(+) + CO3(2-) + H2O=HCO3(—) + OH(—) + 2Na(+) 

CO3(2-) + H2O = HCO3(—) + OH(—)

pH>7


3) NaNO2+ H2O = HNO2 + NaOH 

Na+ + NO2—+ H2O = HNO2 + Na+ + OH— 

NO2—+ H2O = HNO2 + OH—

pH>7


4) KCl+H2О = the reaction does not proceed, since the solution medium is neutral, and pH=7